OpenStudy (anonymous):

A box contains 171 marbles which are either red or blue. There are 25 more blue marbles than red marbles. How many blue marbles are in the box?

hero (hero):

r + b = 171 b = 25 + r Solve the system

OpenStudy (anonymous):

b+r=171 b=r+25 r+25+r=171 2r=146 r=73 b=171-r b=171-73 b=98

OpenStudy (anonymous):

oh that makes so much more sense! so the answer is that there are 98 blue marbles?

OpenStudy (anonymous):

yes